How much time will be needed for $24,000 to grow to $24,974.50 if deposited at 4% compound quaterly?
Answer by Boreal(15235)

24974.50=24000(1+.04/4)^4t
1.041=(1.01)^4t
ln of both sides before rounding.
0.0398=4t(ln1.01)=0.0398
1 year
check
$24000(1.01)^4=$24950.